Page MenuHomePhabricator

2*10G optics down on cr2-esams
Open, HighPublic

Description

Since the upgrade of T243080.

Those two interfaces on cr2-esams are not coming back up:
xe-0/1/4 to Core: cr3-esams:xe-0/0/0 {#20033}
xe-0/1/6 to down Core: cr3-esams:xe-0/1/0 {#20046}

We should try to replace them.

I'm pretty sure we have spare 10G optics on site, but not 100% sure.

We will need remote hands to do the swap.

Event Timeline

ayounsi triaged this task as High priority.Feb 18 2020, 2:53 PM
ayounsi created this task.
Restricted Application added a project: Operations. · View Herald TranscriptFeb 18 2020, 2:53 PM
Restricted Application added a subscriber: Aklapper. · View Herald Transcript

Can probably be combined with T242097.

mark added a subscriber: mark.Feb 18 2020, 2:59 PM

There are multiple 10G LR optics on-site for sure. Longer distance ones, less so.

RobH added a comment.Feb 18 2020, 5:27 PM

There are multiple 10G LR optics on-site for sure. Longer distance ones, less so.

@mark,

Can you advise where those spare optics are being stored for my directions to remote hands? I've looked in the cr[23] and asw2-esams devices, and the only spare optics labeled there are on asw2-esams:

ge-4/0/27       down  down Spare optic SFP-T
ge-5/0/27       down  down Spare optic SFP-LH

I don't see any other spares, please advise.

mark added a comment.Feb 21 2020, 12:17 PM

I am pretty sure there are a bunch of optics (of various kinds) in the "spare" switches, in the bottom of rack OE15. Unfortunately those switches are not powered up, and certainly not configured and remote manageable - something we should probably fix on next visit.

The optics of the right kind can be recognized by the 'color' of their label (probably white, but not 100% sure). Perhaps we can figure it out from one of the pictures of the known ports?

RobH moved this task from Backlog to Break/Fix on the ops-esams board.Feb 21 2020, 4:26 PM

Replacing the optics on both sides didn't help, and light levels are correct.

Service Request ID 2020-0319-0197 has been created.

JTAC's next step is to reboot the FPC... I asked them if there was any other less intrusive ways to test first.